OpenStudy (anonymous):

can anyone help. x/4x+x^2

OpenStudy (anonymous):

what are we suppose to do here?

OpenStudy (anonymous):

solve it, please

OpenStudy (anonymous):

does it equal 1 over 4

OpenStudy (anonymous):

there is no equal sign, how can i solve for anything?

OpenStudy (anonymous):

all you have right now are two terms, do you want me to combine the two terms?

OpenStudy (anonymous):

please, I am stuck

OpenStudy (anonymous):

It say to simplfy them??????

OpenStudy (anonymous):

oh, thats better, now i understand what we need to do

OpenStudy (dumbcow):

im guessing you mean simplify \[\frac{x}{x^{2}+4x} = \frac{x}{x(x+4)} = \frac{1}{x+4}\]

OpenStudy (anonymous):

Oh, so I was close. Thanks so much.